The points A and B have coordinates (-8,-8) and (12,2) respectively AB is the diameter of circle C
(a) find an equation for C
The point ( 4,8) also lies on C
find an equation of the tangent to c at the point (4,8) giving your answer in for of ax + by + c= 0
1
Expert's answer
2020-11-10T09:58:00-0500

(a) The equation of a regular circle with radius R is


"R^2=(x-a)^2+(y-b)^2."

Substitute the point we are given:


"R^2=(-8-a)^2+(-8-b)^2,\\
R^2=(12-a)^2+(2-b)^2."

We have three undefined values. However, we know that the distance between these two given points is the diameter, so, the radius is


"R=\frac D2=\frac{\sqrt{(-8-12)^2+(-8-2)^2}}{2}=5\sqrt5."

Substitute and find a and b:


"125=(-8-a)^2+(-8-b)^2,\\
125=(12-a)^2+(2-b)^2.\\
a=2\\
b=-3"


So, the equation of the circle is


"125=(x-2)^2+(y+3)^2."

(b) To find the equation of a tangent, express y in terms of x and find the derivative:


"y'=(\sqrt{-x^2+4x+121}-3)'=\\\space\\
=\frac{2-x}{\sqrt{-x^2+4x+121}}."

Substitute x-value to find the gradient:


"k=y'(4)=-\frac{2}{11}."

The equation is


"y-y_1=k(x-x_1),\\
y-8=-\frac{2}{11}(x-4),\\\space\\
\frac{2}{11}x+y-\frac{96}{11}=0."